# Predicting from a Different Perspective in Re-ranking Model for Inductive Knowledge Graph Completion

### v1

@@ 5.1 Model details
In this section, we introduce ReDistLP, a re-ranking model for predicting missing links. Figure 2 illustrates an overview of our model. ReDistLP is a cascaded model that first scores candidate entities using an initial retriever, and then re-ranks them with a re-ranker. As discussed in Section 3, we aim to select the initial retriever and re-ranker models such that their prediction sets ($S_{m_{1}}$ and $S_{m_{2}}$) satisfy the condition ($S_{m_{1}}\not\subset S_{m_{2}}$), according to equation (7). This non-subset relationship between the prediction sets can enhance the effectiveness of the re-ranking process. As shown in the analysis in Section 4, when trained with longer hop rules, the BERTRL tends to predict entities concentrated within narrower regions of the KG. To satisfy the condition in equation (7) and maximize re-ranking effectiveness, we employ a BERTRL trained on shorter hop rules as the initial retriever, and a BERTRL trained on longer hop rules as the re-ranker Therefore, in our final ReDistLP, we use BERTRL trained with 2-hop as the initial retriever and BERTRL trained with 3-hop as the re-ranker.
@@
We make two variants of datasets for BERTRL in training and inference. The first is the same as the original one. This variants dataset excludes the triples where the head and tail entities do not connect within $n$-hop. The second variant does not exclude such triple data. Table 3 shows examples of training data. The first dataset only contains No.1 and 2 in this case, while the second dataset contains all cases. This enables the model to score entities that do not connect to the other entity in a triple. We call BERTRL tested on the second dataset as BERTRL v2. We use BERTRL v2 as the initial retriever for our model.
@@
5.2 Rank Threshold Optimization
Determining the optimal cutoff point for the initial retriever’s results directly affects the accuracy of the re-ranking model. Ideally, the cutoff would be defined as the position of the answer entity. Actually, in practice, model predictions are continuous probability scores, which can be noisy, making it difficult to identify a stable threshold. We explored different threshold optimization methods to find the most suitable approach for our model.
Constant Threshold One straightforward approach is to use a constant threshold, where a fixed threshold is applied to the ranking list. We determine the optimal fixed threshold using a development dataset and then apply it to the test set.
Top-$k$ Cutoff Another common approach is to use a top-k cutoff, where the top-$k$ items from the ranked list are selected. This method ensures a fixed number of items are always selected, regardless of their absolute scores.
k-means clustering A third approach is to use k-means clustering, where the items in the ranked list are grouped into k clusters based on their similarity. Items belonging to the top-scoring cluster(s) are selected. This method groups similar items together and selects the most relevant cluster(s) based on their overall scores.
Continuous approach The discrete methods mentioned above have a major drawback: they often exclude the answer entity from the selected set. When the answer entity is not included in the initial pool, the accuracy of the re-ranking model is limited to that of the initial retriever. In contrast, a continuous approach can be employed to mitigate this issue. Instead of applying fixed cutoff points, this method passes the results with their predicted scores to the re-ranker. The re-ranker then re-ranks the results while consulting the scores of the initial retriever. This approach allows for a more flexible selection process.
  • Mean
    This method calculates the final score by simply averaging the prediction scores from the two models. Our final design uses this method.
  • Mean of Normalized Scores (MNS)
    The confidence for a specific input may differ between the two models. To account for this, we normalize each score before averaging them.
  • Reciprocal Rank (RR)
    In this method, we calculate the reciprocal rank of each entity in the prediction sets of both models. The final score is determined by the mean of these reciprocal ranks.
  • Reciprocal Rank with Score (RRS)
    In this method, we multiply the score by the reciprocal rank for each entity in both models. The final score is determined by the mean of these values.
@@ 6 Experiments
In our experiments, we evaluate the effectiveness of our proposed model on inductive knowledge graph completion datasets (Section 6.4). In the latter part of Section 6.4, we analyze the performance of our model in the ideal setting. Subsequently, we conduct an ablation study in Section 6.5.
6.1 Datasets
We use inductive subsets of three knowledge completion datasets: WN18RR (Dettmers et al., 2018), FB15k-237 (Toutanova et al., 2015), and NELL-995 (Xiong et al., 2017), which were introduced by Teru et al. (2020). WN18RR is a subset of WordNet that stores lexical relations of words, with its inverse relations removed. FB15k-237 is a subset of Freebase that stores general knowledge facts. NELL-995 is a subset of NELL that is constructed from the web automatically. Table 4 provides a detailed overview of the statistics of KG datasets.
Table 4: Number of entities, relations, and triples in training and inference graphs.
Dataset Graph $||\mathcal{E}|$$||\mathcal{R}|$$|\mathcal{T}|$
WN18RR train 2746 9 6670
train-2000 1970 9 2002
train-1000 1362 9 1001
test-inductive 922 8 1991
FB15k-237 train 1594 180 5223
train-2000 1280 180 2008
train-1000 923 180 1027
test-inductive 1093 142 2404
NELL-995 train 2564 88 10063
train-2000 1346 88 2011
train-1000 893 88 1020
test-inductive 2086 79 6621
6.2 Baselines and Implementation Details
We compare our approach with GraIL (Teru et al., 2020), BERTRL (Zha et al., 2022), KG-BERT (Yao et al., 2019), RuleN (Meilicke et al., 2018), and KRST (Su et al., 2023). GraIL is a GNN-based model that predicts over subgraph structures. KRST is a transformer-based rule-induction method that selectively incorporates reliable rules during training. We use the implementation provided publicly by the authors and the best hyper-parameter settings reported in their work.
@@ 6.3 Experiment Settings
We report the evaluation metrics Hits@1 and Mean Reciprocal Rank (MRR). Hits@1 measures the percentage of cases where the correct triple is ranked as the top 1 prediction. MRR computes the average of the reciprocal ranks of the correct triple. Consistent with the evaluation setup in Teru et al. (2020) and other prior work, we measure the rank of the correct triple among a set of 50 candidate triples, comprising the correct triple and 49 negative triples.
@@ 6.4 Results and Analysis
Actual Setting. Table 5 and Table 6 present the Hits@1 and MRR results of inductive link prediction, respectively. We adopt the BERTRL v2 with 2-hop as the initial retriever and the BERTRL with 3-hop as the re-ranker of our final model.
Table 5: Hits@1 of link prediction in the inductive setting.
@@
RuleN .745 .415 .638
GRAIL .769 .390 .554
KG-BERT .436 .341 .244
BERTRL .755 .541 .715
KRST .835 .602 .649
BERTRL (1 hop) .630 .329 .466
BERTRL (2 hop) .729 .410 .639
BERTRL (3 hop) .793 .578 .731
ReDistLP (2 to 3-hop) .796 .634 .748
Table 6: MRR of link prediction in the inductive setting.
WN18RR FB15k-237 NELL-995
RuleN .780 .462 .710
GRAIL .799 .469 .675
KG-BERT .542 .500 .419
BERTRL .792 .605 .808
KRST .902 .716 .769
BERTRL (1 hop) .630 .344 .466
BERTRL (2 hop) .726 .415 .637
BERTRL (3 hop) .718 .595 .726
ReDistLP (2 to 3-hop) .808 .715 .832
Our model achieves the best performance under Hits@1 on the FB15k-237 and NELL-995 datasets. Specifically, ReDistLP achieves significant improvements in the FB15k-237 (+3.2% for Hits@1) and NELL-995 (+3.3% for Hits@1). Under the MRR metric, our model achieves the best performance on the NELL-995 dataset and comparable results on the FB15k-237 dataset.
Ideal Setting. Table 7 presents the results of inductive link prediction in the ideal setting. In this ideal setting, we determine the cutoff based on the position of the answer entity in the initial pool. Using this optimal cutoff, we adopt the top-$k$ re-ranking algorithm.
@@
Hits@1
ReDistLP (1 to 2-hop) .725 .590 .745
ReDistLP (1 to 3-hop) .784 .720 .827
ReDistLP (2 to 3-hop) .817 .705 .831
MRR
ReDistLP (1 to 2-hop) .742 .677 .818
ReDistLP (1 to 3-hop) .806 .781 .891
ReDistLP (2 to 3-hop) .836 .777 .893
The Hits@1 metrics show a large difference between the actual setting and the ideal setting. Specifically, Hits@1 of ReDistLP (2 to 3-hop) in the ideal setting is +2.6%, +7.1%, +8.3% higher than in the actual setting on the WN18RR, FB15k-237, NELL-995, respectively. This indicates if we could determine a more optimal cutoff, our model’s accuracy could be improved. In Section 6.5, we present an ablation study on the selection of the ranking strategy.
ReDistLP (2 to 3-hop) shows the best results. However, when compared with the initial retriever of BERTRL (Table 5, 6), ReDistLP (1 to 3-hop) demonstrates the largest improvements. To better understand the factors contributing to these differences in performance, we investigate the difference between their prediction sets. We calculate the following equation
$r=\frac{|S_{G}\setminus S_{R}|}{|S_{G}|},$ (9)
where $S_{G}$ and $S_{R}$ represent the prediction sets of the initial retriever and the re-ranker, respectively. This equation, which corresponds to equation (8), normalizes the difference in cardinality between the two prediction sets. A higher value of $r$ indicates a larger difference between the two sets. Table 8 shows the mean of $r$ calculated over the test set.
@@
6.5 Ablastion study
Table 9 shows the accuracy of different rank threshold optimization algorithms in ReDistLP. The averaging method achieves the best results among them. In the NELL-995 dataset, the RRS method outperforms the Mean method.
Overall, the discrete methods, such as k-means, top-k, and fixed threshold, could not outperform the continuous methods. However, from Table 7, even discrete methods have the potential to improve accuracy. This implies that determining the optimal cutoff in the initial retriever’s results is challenging in our model, as the set must be minimized while still including the correct entity. We will address this issue in the future study.
Table 9: Ablation study of ReDistLP variants (Hits@1)
WN18RR FB15k-237 NELL-995
top-$k$ cutoff ($k=5$) .737 .620 .702
top-$k$ cutoff ($k=10$) .739 .615 .708
top-$k$ cutoff ($k=20$) .749 .617 .722
top-$k$ cutoff ($k=40$) .780 .607 .736
fixed thresholds ($\theta=0.5$) .763 .615 .731
fixed thresholds ($\theta$ determined by the dev set) .794 .627 .735
k-means (1 out of 2) .775 .600 .704
k-means (3 out of 5) .775 .620 .719
Mean .796 .634 .748
Max .761 .627 .746
Min 0.794 0.600 0.721
MNS .778 .620 .741
RR .749 .593 .723
RRS .778 .627 .751
@@ 7 Conclusion
We proposed ReDistLP, a re-ranking model designed for inductive link prediction on KGs. ReDistLP employs two variants of BERTRL, which are trained with different rules. Different rules change the model’s behavior. Specifically, the model trained with longer hop rules predicts entities within narrower regions of the KG. ReDistLP leverages this phenomenon through a re-ranking pipeline to improve link prediction performance. Our model outperforms the baselines in 2 out of 3 datasets.
Furthermore, we demonstrated that our model’s performance can be further improved if an optimal rank threshold for re-ranking could be determined. In the future, we plan to optimize the rank threshold for our model.

### v2

@@ 4.2 Model Details
Figure 2 provides an overview of our ReDistLP model. ReDistLP is a cascaded model that first scores candidate entities using an initial retriever and then re-ranks them with a re-ranker. As discussed in Section 4.1, we aim to select the initial retriever and re-ranker models such that their prediction sets ($S_{m_{1}}$ and $S_{m_{2}}$) satisfy the condition ($S_{m_{1}}\not\subset S_{m_{2}}$), according to equation (8). This non-subset relationship between the prediction sets can enhance the effectiveness of the re-ranking process. As shown in the analysis in Section 3, when trained with longer hop rules, the BERTRL tends to predict entities concentrated within narrower regions of the KG. To satisfy the condition in equation (8) and maximize re-ranking effectiveness, we employ a BERTRL trained on shorter hop rules as the initial retriever and a BERTRL trained on longer hop rules as the re-ranker Therefore, in our final ReDistLP, we use BERTRL trained with 2-hop as the initial retriever and BERTRL trained with 3-hop as the re-ranker.
The initial retriever scores candidate entities $C$ using BERTRL trained with 2-hop rules. The prediction of the initial retriever is as follows:
$S_{I}(h,r)=\bigl{\{}((h,r,t),score_{{\text{BERTRL 2-hop}}}(h,r,t))|t\in C\bigr{\}}$ (10)
The re-ranker scores the candidate entities using BERTRL trained with 3-hop rules, denoted as $S_{R}(h,r)$, which follows a similar definition to equation (10). The re-ranked prediction set $S^{\prime}(h,r)$ is obtained from the initial prediction set $S_{I}(h,r)$ and the scores $S_{R}(h,r)$, where the final score for each entity $e$ is given by $\mu_{S^{\prime}(h,r)}(e)$.
4.3 Rank Threshold Optimization
The optimal cutoff point for the initial retriever’s results directly affects the accuracy of the re-ranking model. Ideally, the cutoff would be defined as the position of the answer entity. However, in practice, model predictions are continuous probability scores, which can be noisy, making it difficult to identify a stable threshold. We explored different threshold optimization methods to find the most suitable approach for our model.
4.3.1 Classical Sets Approach.
Classical sets, also known as crisp sets, have well-defined boundaries where an element either belongs to the set or does not. In the context of re-ranking, we can apply classical set theory to determine where to truncate the ranking list generated by the initial retriever.
Threshold.
One straightforward approach is to use a constant threshold, where a fixed threshold is applied to the ranking list. The re-ranking function using a threshold $\theta$ is defined as
$\operatorname{re-rank}_{\theta}(S_{I},S_{R})=\Bigl{\{}\Bigl{(}(h,r,t),\mu_{S_{R}}(t)\Bigr{)}|t\in S_{I}(h,r)^{\geq\theta}\Bigr{\}}.$ (11)
The re-ranking function selects the triples from the $\theta$-cut of the initial prediction set. We determine the optimal fixed threshold $\theta$ using a development dataset and then apply it to the test set.
Top-$k$ Cutoff.
Another common approach is to use a top-k cutoff, where the top-$k$ items from the ranked list are selected. This method ensures a fixed number of items are always selected, regardless of their absolute scores. The re-ranking function using a top-$k$ cutoff is given by:
$\operatorname{re-rank}_{{\text{t}op}\mbox{-}k}(S_{I},S_{R})=\Bigl{\{}\Bigl{(}(h,r,t),\mu_{S_{R}}(t)\Bigr{)}|t\in{\text{t}op}\mbox{-}k(S_{I}(h,r))\Bigr{\}}$ (12)
where the function ${\text{t}op}\mbox{-}k(S_{I}(h,r))$ selects the top-$k$ entities from the initial prediction set based on their membership scores.
$k$-means clustering.
A third approach is to use k-means clustering, where the items in the ranked list are grouped into $k$ clusters based on their similarity. Items belonging to the top-scoring cluster(s) are selected for re-ranking. This method groups similar items together and selects the most relevant cluster(s) based on their overall scores. The re-ranking function using $k$-means clustering can be expressed as:
$\operatorname{re-rank}_{k\mbox{-}{\text{m}eans}}(S_{I},S_{R})=\Bigl{\{}\Bigl{(}(h,r,t),\mu_{S_{R}}(t)\Bigr{)}|(h,r,t)\in\bigcup_{j=1}^{m}C_{j}\Bigr{\}}$ (13)
where the function $\bigcup_{j=1}^{m}C_{j}$ represents the union of the top-$m$ clusters obtained from applying $k$-means clustering to the probability distribution of the initial prediction set $S_{I}(h,r)$.
4.3.2 Fuzzy Sets Approach.
The classical sets approach has a major drawback: it often excludes the answer entity from the selected set. When the answer entity is not included in the initial pool, the re-ranking model’s accuracy is limited to that of the initial retriever. In contrast, the fuzzy sets approach can be employed to mitigate this issue.
Intersection.
The equation (5) is equal to $S_{m_{1}}^{\geq\mu_{m_{1}}(\bar{t})}\cap S_{m_{2}}^{\geq\mu_{m_{2}}}(\bar{t})$. Generalizing this approach leads us to the definition of the intersection-based re-ranking function: $\operatorname{re-rank}_{\cap}(S_{I},S_{R})=S_{I}\cap S_{R}$. The membership function of the resulting fuzzy set is given by:
$\mu_{\operatorname{re-rank}_{\cap}(S_{i},S_{r})}(t)={\text{m}in}(\mu_{S_{I}}(t),\mu_{S_{R}}(t))$ (14)
Union.
In fuzzy set theory, the union of two fuzzy sets can be defined using the maximum operator (Zadeh, 1965). Applying this concept, we introduce the union-based re-ranking function $\operatorname{re-rank}_{\cup}(S_{I},S_{R})$. The membership function of the resulting fuzzy set for the union-based re-ranking approach is given by:
$\mu_{\operatorname{re-rank}_{\cup}(S_{i},S_{r})}(t)={\text{m}ax}(\mu_{S_{I}}(t),\mu_{S_{R}}(t))$ (15)
Convex combination (Mean).
Another approach to combining fuzzy sets is through a convex combination. The convex combination of $A$, $B$, and $\Lambda$ is represented by $(A,B;\lambda)$ (Zadeh, 1965). The intersection, convex combination, and union of fuzzy sets satisfy the following relationship (Zadeh, 1965):
$A\cap B\subset(A,B;\Lambda)\subset A\cup B\quad\text{for all $\Lambda$}$ (16)
This relationship implies that the convex combination of two fuzzy sets is always bounded by their intersection and union, regardless of the weight fuzzy set $\Lambda$ choice. The convex combination has a softened membership function compared with the intersection. Hence, we introduce the mean-based re-ranking function $\operatorname{re-rank}_{\text{mean}}(S_{I},S_{R})$ as follows:
$\operatorname{re-rank}_{{\text{m}ean}}=(S_{I},S_{R};\Lambda)$ (17)
where $\Lambda$ is a fuzzy set representing the weights of the convex combination. We define $\Lambda$ as a fuzzy set with a membership value of 0.5 for all entities. The membership function of the resulting fuzzy set is given by:
$\mu_{(S_{I},S_{R};\Lambda)}(t)=\mu_{\Lambda}(t)\mu_{S_{I}}(t)+(1-\mu_{\Lambda}(t))\mu_{S_{R}}(t)$ (18)
Our final model uses this re-ranking function.
4.4 Training
@@
We make two variants of datasets for BERTRL in training and inference. As the original one used in the previous study, the first variant of the dataset excludes the triples where the head and tail entities do not connect within $n$-hop. The second variant does not exclude such triple data. Table 3 shows examples of training data. In this case, the first dataset only contains No.1 and 2, while the second dataset contains all cases. This enables the model to score entities that do not have any connecting path to the other entity in a given triple within the n-hop. We call BERTRL tested on the second dataset as BERTRL v2, which is used as the initial retriever for our model.
@@ 5 Experiments
In our experiments, we evaluate the effectiveness of our proposed model on inductive knowledge graph completion datasets. In the latter part, we analyze the performance of our model in the ideal setting.
@@ 5.1 Experiment Settings
We use inductive subsets of three knowledge completion datasets: WN18RR (Dettmers et al., ), FB15k-237 (Toutanova et al., ), and NELL-995 (Xiong et al., ), which were introduced by Teru et al. (Teru et al., ).
We report the evaluation metrics Hits@1 and MRR. Hits@1 measures the percentage of cases where the correct triple is ranked as the top 1 prediction. MRR computes the average of the reciprocal ranks of the correct triple. Consistent with the evaluation setup in GRAIL (Teru et al., ) and other prior work, we measure the rank of the correct triple among a set of 50 candidate triples, comprising the correct triple and 49 negative triples.
We compare our approach with GraIL (Teru et al., ), BERTRL (Zha et al., ), KG-BERT (Yao et al., 2019), RuleN (Meilicke et al., ), and KRST (Su et al., ). We use the implementation provided publicly by the authors and the best hyper-parameter settings reported in their work.
@@ 5.2 Results and Analysis
Actual Setting. Table 4 presents the Hits@1 and MRR results of inductive link prediction.
Table 4: Hits@1 and MRR of link prediction in the inductive setting.
@@
Hits@1 MRR Hits@1 MRR Hits@1 MRR
RuleN .745 .780 .415 .462 .638 .710
GRAIL .769 .799 .390 .469 .554 .675
KG-BERT .436 .542 .341 .500 .244 .419
BERTRL .755 .792 .541 .605 .715 .808
KRST .835 .902 .602 .716 .649 .769
BERTRL (1 hop) .630 .661 .329 .390 .466 .519
BERTRL (2 hop) .729 .749 .410 .455 .639 .698
BERTRL (3 hop) .793 .815 .578 .636 .731 .821
ReDistLP (2 to 3-hop) threshold $\theta$794 - .627 - .735 -
ReDistLP (2 to 3-hop) top-$k$ cutoff ($k=40$) .780 - .607 - .736 -
ReDistLP (2 to 3-hop) $k$-means (3 out of 5) .775 - .620 - .719 -
ReDistLP (2 to 3-hop) Intersection .794 .814 .600 .631 .721 .805
ReDistLP (2 to 3-hop) Union .761 .793 .627 .706 .746 .831
ReDistLP (2 to 3-hop) Mean .796 .808 .634 .715 .748 .832
Our model achieves the best performance under Hits@1 on the FB15k-237 and NELL-995 datasets. Specifically, ReDistLP achieves significant improvements in the FB15k-237 (+3.2% for Hits@1) and NELL-995 (+3.3% for Hits@1). Under the MRR metric, our model achieves the best performance on the NELL-995 dataset and comparable results on the FB15k-237 dataset. In WN18RR, however, KRST obtains the highest. Although our model does not achieve the best performance on WN18RR, it still outperforms BERTRL under the Hits@1 and MRR metrics. ReDistLP, which combines a 2-hop retriever and a 3-hop re-ranker, outperforms BERTRL (3-hop) on all datasets for both Hits@1 and MRR metrics, demonstrating the effectiveness of our redistribution approach.
Ideal Setting. Table 5 presents the results of inductive link prediction in the ideal setting. In this ideal setting, we determine the cutoff based on the position of the answer entity in the initial pool.
@@
H@1 MRR H@1 MRR H@1 MRR Avg. MRR
ReDistLP (1 to 2-hop) top-$k$725 .742 .590 .677 .745 .818 .746
ReDistLP (1 to 3-hop) top-$k$784 .806 .720 .781 .827 .891 .826
ReDistLP (2 to 3-hop) top-$k$817 .836 .705 .777 .831 .893 .835
ReDistLP (2 to 3-hop) Intersection .820 .850 .695 .781 .742 .827 .819
ReDistLP (2 to 3-hop) Union .775 .797 .671 .738 .805 .872 .802
ReDistLP (2 to 3-hop) Mean .801 .821 .661 .736 .772 .850 .802
The Hits@1 metrics show a large difference between the actual setting and the ideal setting. Specifically, the Hits@1 of ReDistLP (2 to 3-hop) in the ideal setting is higher than in the actual setting by 2.6%, 7.1%, and 8.3% on WN18RR, FB15k-237, and NELL-995, respectively. This indicates if we could determine a more optimal cutoff, our model’s accuracy could be improved.
ReDistLP (2 to 3-hop) shows the best results. However, when compared with the initial retriever of BERTRL (Table 4), ReDistLP (1 to 3-hop) demonstrates the largest improvements. To better understand the factors contributing to these differences in performance, we investigate the difference between their prediction sets. We calculate the following equation
$r=\frac{|S_{I}\setminus S_{R}|}{|S_{I}|},$ (19)
where $S_{I}$ and $S_{R}$ represent the prediction sets of the initial retriever and the re-ranker, respectively. This equation, which corresponds to equation (9), normalizes the difference in cardinality between the two prediction sets. A higher value of $r$ indicates a larger difference between the two sets. Table 6 shows the mean of $r$ calculated over the test set.
@@
Compared to ReDistLP variants, the classical set approaches, such as $k$-means, top-$k$, and fixed threshold, could not outperform the fuzzy set approach in the actual setting. In the ideal setting, however, the classical approach achieves better Hits@1 and Avg. MRR results than all of the fuzzy set approaches. This implies that determining the optimal cutoff in the initial retriever’s results is challenging in our model. The set must be minimized while still ensuring that the correct entity is included. We will address this issue in future studies to improve its performance further.
@@ 6 Conclusion
We proposed ReDistLP, a re-ranking model designed for inductive link prediction on KGs. ReDistLP employs two variants of BERTRL, which are trained with different rules. We analyzed that different rules lead to changes in the model’s behavior. Specifically, the model trained with longer hop rules predicts entities within narrower regions of the KG. ReDistLP leverages this phenomenon through a re-ranking pipeline to improve link prediction performance. Our model outperforms the baselines in 2 out of 3 datasets.
Furthermore, we showed that our model’s performance can be further improved if an optimal rank threshold for re-ranking could be determined. In the future, we plan to optimize the rank threshold for our model.